OpenStudy (anonymous):

Kathie ran for twice as long as she walked. She walked for 10 minutes more than she rode her bike. She walked for 30 minutes. How many minutes did Kathie exercise altogether? Choose all answers that are correct. A. Add 30 + 30. Subtract 10 from 30. Add the sum and the difference. Add that sum to 30. B. Multiply 30 × 2. Subtract 10 from 30. Add the product and the difference. Add that sum to 30. C. Divide 30 by 2. Then add the quotient to 10. D. Multiply 30 × 2. Add that product to 30. Divide the sum by

OpenStudy (reemii):

Give a name to the durations "Run", "Walk", "ride Bike" : \(R,W,B\). You are told \(W=30\). - What is \(R\)? - What is \(B\)? - What is total exercise time?

OpenStudy (reemii):

Translate the statement into maths: `Kathie ran for twice as long as she walked.` -> \(R = 2W\).

OpenStudy (anonymous):

So, it would be, 2 x 30?

OpenStudy (reemii):

Yes.

OpenStudy (reemii):

How do you translate `She walked for 10 minutes more than she rode her bike.` into maths ?

OpenStudy (anonymous):

60

OpenStudy (anonymous):

2 x 30 = 60

OpenStudy (anonymous):

Just 10 x ?

OpenStudy (anonymous):

Adoy.

OpenStudy (anonymous):

I meant 10 x 60

OpenStudy (reemii):

`more` will translate to something an expression with the operation "+" or "-" .

OpenStudy (anonymous):

Lemme write that down. Okay, got it.

OpenStudy (anonymous):

@reemii

OpenStudy (anonymous):

Are you there?

OpenStudy (reemii):

I am

OpenStudy (anonymous):

I kinda lost connection there.

OpenStudy (anonymous):

Fck this test is due half a hour.

OpenStudy (reemii):

\(X\) is ten more than \(Y\) translates to \(X=Y+10\). How do you translate `She walked for 10 minutes more than she rode her bike.` in terms of \(R\) and \(B\)? `She walked for 10 minutes more than she rode her bike.`= `Walk is ten more than Bike`. etc. next step is to write the equation.

OpenStudy (reemii):

that will be \(W = B+10\). So, \(B=W-10=30-10\). The total amount of time is \(W+B+R\).

OpenStudy (anonymous):

Okay, I got the rest thanks!